You're rolling solid rubber balls on the kitchen floor. Ball 1 has a density of 1.00 ×

×10

3

3kg/m

3

kg/m3and a radius of 30.0 mm

mm . Ball 2 has an unknown density and a radius of 43.0 mm

mm and is initially at rest. You roll ball 1 at an initial speed of 3.00 m/s

m/s, and the two balls collide head-on. Ball 1 reverses direction and comes back to you at 2.00 m/s

m/s, and after the collision, the speed of ball 2 is 1.00 m/s

m/s. The positive x

x axis is in the direction of ball 1's initial motion. 1.What is the initial momentum of ball 2.

2 what is the density of ball 2

3. What is the initial momentum of ball 1



1
Expert's answer
2021-04-05T11:10:24-0400

3)


"p_1=\frac{4}{3}\pi(0.03)^3(1000)(3)=0.339\frac{kgm}{s}"

2)


"\frac{4}{3}\pi(0.03)^3(1000)(3+2)=\frac{4}{3}\pi(0.043)^3(1)\rho_2\\\rho_2=1700\frac{kg}{m^3}"

1)


"p_2=0\frac{kgm}{s}"